OpenStudy (anonymous):

Suppose that 53% of us sleep between 7 and 8 hours a day. Express this percent as a decimal.

OpenStudy (anonymous):

x% = x/100 So for example, 23% = 23/100 23% = 23/100 = 0.23 which means 23% = 0.23

OpenStudy (anonymous):

where did 23% come from

OpenStudy (anonymous):

I'm giving am example. Substitute the numbers with the ones in your problem to find the answer.

OpenStudy (anonymous):

*an

OpenStudy (anonymous):

ok thanks !

OpenStudy (anonymous):

I thought I had to do something with the 7 and 8

OpenStudy (anonymous):

No, that's how they wanted to trick you, to make it seem more difficult than it is.The question states: Express this [percent] as a decimal. You only had to work with the percent.

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

You're Welcome, and Welcome to Open Study!!!:D Let me know if you have any questions with anything:)